. From a circular sheet of radius 4 cm, a circle of radius 3 cm is removed. Find the area of the remaining sheet. (Take π = 3.14)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is give that,
Radius of circular sheet R = 4 cm
A circle of radius to be removed r = 3 cm
Then,
The area of the remaining sheet =π R^2-π r^2
=π (R^2–r^2)
= 3.14 (42- 32)
= 3.14 (16 - 9) = 3.14 × 7
= 21.98 cm^2
So, the area of the remaining sheet is 21.98 cm^2

Thanks